# Constant-Time Motion Planning with Manipulation Behaviors

### v1

@@ VI Conclusion and Future Work
This work introduces B-CTMP, a constant-time algorithm that extends existing CTMP methods to incorporate manipulation behaviors directly into the preprocessing phase. Instead of treating collision-free motion planning and behavior execution as separate sequential processes, B-CTMP integrates behavior validation during offline computation to improve the reliability of manipulation tasks in semi-structured environments. The method addresses a practical limitation in current constant-time planning approaches: the tendency to plan paths to kinematically feasible states that may prove unsuitable for successful behavior execution. A key insight is that by defining the region of interest in object space rather than robot configuration or task space, B-CTMP can automatically discover the relevant initiation states for the manipulation task. This eliminates the need for manual specification of robot configuration or task-space initiation regions, which requires human intuition and domain expertise. Through behavior validation during preprocessing and selection of initiation states based on their coverage properties, B-CTMP reduces the likelihood of execution failures that commonly occur when motion planning and behavior execution are decoupled. The experimental results demonstrate that B-CTMP maintains consistent success rates while achieving fast online query performance. Baseline methods exhibit higher failure rates, particularly in challenging regions where kinematically reachable states prove behaviorally invalid. The approach also demonstrates improved memory efficiency by organizing neighborhoods in object space. We introduce PR-completeness as a framework for reasoning about solution guarantees in behavior-based manipulation tasks within predefined regions of interest. Several directions for future work emerge from this work. The approach could be extended to handle sequences of multiple behaviors. The preprocessing framework may also provide a foundation for learning-based methods where behavior policies are refined or analyzed during offline computation to increase the probability of successful execution online. B-CTMP provides a practical step toward more reliable preprocessing-based manipulation planning by demonstrating that behavior validation can be effectively integrated into constant-time frameworks. While the approach focuses on specific classes of two-step manipulation tasks, it offers a useful building block for developing more robust robotic automation systems that can provide predictable performance.

### v2

@@ VII Conclusion
B-CTMP introduces a constant-time algorithm that integrates manipulation behaviors directly into the preprocessing phase, moving beyond methods that treat motion planning and behavior execution as decoupled, sequential processes. By performing behavior validation during offline computation, the method addresses a core limitation in current constant-time planning: the tendency to reach kinematically feasible states that prove unsuitable for task execution. A key to our approach is the definition of the region-of-interest in object space rather than configuration or task space. This approach automatically discovers relevant initiation states, eliminating the need for manual specification or human domain expertise. Experimental results demonstrate that B-CTMP achieves consistent success rates and fast online performance, while baseline methods fail. Finally, we introduce PR-completeness as a framework for reasoning about solution guarantees in behavior-based manipulation tasks. Future work could extend this approach to handle sequences of multiple behaviors or provide a foundation for learning-based methods where policies are refined during offline computation. Ultimately, B-CTMP provides a robust building block for predictable, behavior-aware robotic automation in semi-structured environments.